I have a 10 foot in diameter fountain if I build a 3 foot walkway around it what would be the square footage of the walkway?

I would like the answer and how to get the answer thank you so much zina

Answer by jim_thompson5910(35256)

Area of fountain only:

A = pi*r^2

A = pi*5^2

A = 25pi

----------------------
Area of fountain+walkway

A = pi*r^2

A = pi*8^2

A = 64pi
---------------------
Area of walkway only

A = (Area of fountain+walkway) - (Area of fountain)

A = 64pi - 25pi

A = 39pi

So the area of the walkway only is exactly 39pi square feet. This is approximately 39pi = 39*3.14159 = 122.52201 square feet